Hi, I'm having a lot of trouble figuring out how to do these questions. PLEASE HELP!!!

Question 1. A Solution X is made by taking 10.00 mL of a 1.0 M solution and add enough water to make 100.0 mL total volume. Solution Y is made by taking 5.0 mL of solution X and diluting it again by adding it to 95 mL of water.

a) What is the molarity of Solution X?

b)What is the molarity of solution Y?

Question 2. Calculate the molarity of the solution created when 50.0 mL of 0.132 M NaOH is mixed with 75.0 mL of 0.120 M NaOH (assuming that volumes are additive).

Question 3. If a solution has a measured concentration of OH- of 1.20x10-4 M, what is the pH of this solution?
1
Expert's answer
2018-03-07T12:01:08-0500
Question 1. Molarity = moles of solute/liters of solution
a) n = C*V = 1mol/L*0.01L = 0.01 mol
Cx = 0.01mol/0.1L = 0.1 mol/L
b) n = C*V = 0.1mol/L*0.005L = 0.0005 mol
Cy = 0.0005mol/0.1L = 0.005 mol/L
Question 2. n1(NaOH)=0.132mol/L*0.05L=0.0066 mol
n2(NaOH)=0.120mol/L*0.075L=0.009 mol
C(NaOH)=(0.0066+0.009)mol/(0.05+0.075)L=0.0156mol/0.125L=0.1248 mol/L
Question 3.
[OH-]=1.2*10-4 M
pH + pOH = 14
pH = 14 – pOH
pOH = -lg[OH-] = -lg[1.2*10-4] = -lg[0.00012] = 3.92
pH = 14 - 3.92=10.08